# Early Timestep Zero-Shot Candidate Selection for Instruction-Guided Image Editing

### v1

@@ Abstract
Despite recent advances in diffusion models, achieving reliable image generation and editing results remains challenging due to the inherent diversity induced by stochastic noise in the sampling process. Particularly, instruction-guided image editing with diffusion models offers user-friendly editing capabilities, yet editing failures, such as background distortion, frequently occur across different attempts. Users often resort to trial and error, adjusting seeds or prompts to achieve satisfactory results, which is inefficient. While seed selection methods exist for Text-to-Image (T2I) generation, they depend on external verifiers, limiting their applicability, and evaluating multiple seeds increases computational complexity, reducing practicality. To address this, we first establish a new multiple-seed-based image editing baseline using background consistency scores, achieving Best-of-N performance without supervision. Building on this, we introduce ELECT (Early-timestep Latent Evaluation for Candidate selecTion), a zero-shot framework that selects reliable seeds by estimating background mismatches at early diffusion timesteps, identfying the seed that retains the background while modifying only the foreground. ELECT ranks seed candidates by a background inconsistency score, filtering unsuitable samples early based on background consistency while fully preserving editability. Beyond standalone seed selection, ELECT integrates into instruction-guided editing pipelines and extends to Multimodal Large-Language Models (MLLMs) for joint seed + prompt selection, further improving results when seed selection alone is insufficient. Experiments show that ELECT reduces computational costs (by 41% on average and up to 61%) while improving background consistency and instruction adherence, achieving around 40% success rates in previously failed cases—without any external supervision or training.
Keywords Instruction-guided Image Editing  $\cdot$ Diffusion Models  $\cdot$ Test-Time Scaling  $\cdot$ Seed Selection
Figure 1: Instruction-guided image editing models are highly influenced by the noise seed. To address this issue, we propose a unique candidate selection method (ELECT), which is successfully selects the best seed for background consistency while maintaining the editability of the base model. We propose Background Inconsistency Score ($S^{\text{BIS}}$) that quantifies the degree of unintended background changes in an edited image, measuring relatively how well the background is preserved compared to other candidates in a self-supervised manner. Samples with low $S^{\text{BIS}}$ (blue) have a consistent background while samples with high $S^{\text{BIS}}$ (red) display multiple artifacts and distortion.
@@ 1 Introduction
This inefficiency in manual selection has parallels with inference-time scaling strategies in auto-regressive models like LLMs [38], where multiple generations are sampled to improve output quality. Similarly, in text-to-image generation, search techniques such as Best of N, which select the best result using specific verifiers, have been explored [28]. In instruction-guided image editing, users often generate multiple outputs by varying the random seed and manually selecting the most suitable result. However, this process is computationally expensive, with performance scaling linearly with N, making it impractical for real-time applications. Moreover, there is no established framework for efficiently identifying the optimal seed before full inference, underscoring the need for a more effective selection method.
@@
  • We extend ELECT to joint seed and prompt selection using multimodal large language models (MLLMs) [34] refining out-of-distribution instructions and improving VIEScore by $+0.56$ on average.
@@ 2 Related Work
Instruction-Guided Image Editing with Diffusion Models.
Unlike caption-based approaches, instruction-guided editing methods [1, 50, 9, 37, 44, 52, 7, 16, 18, 8, 48, 24, 53] take an input image $I$ and a textual command $T$ (e.g., ”add a dog”) to guide modifications. InstructPix2Pix (IP2P) [1] uses GPT-3 [2] and Prompt2Prompt [13] to create a dataset, and trains a denoising network conditioned on edit instructions and the original image. Subsequent studies such as MagicBrush [50], UltraEdit [53], HIVE [52], and HQ-Edit [18] propose fine-tuning techniques for IP2P through improved datasets, verifier models [49], and RLHF [26]. InstructDiffusion (InsDiff) [9] proposes a unified framework across multiple computer vision tasks including segmentation and editing, and MGIE [7] and SmartEdit [16] leverages multi-modal LLMs (MLLMs) to guide and enhance image editing. Although instruction-guided methods outperform previous methods, they tend to overedit images and introduce variability due to sensitivity to initial seeds and instruction phrasing, leading to inconsistent edits across runs. To address this, Watch Your Steps [31], Focus on Your Instruction [10], and ZONE [25] employ mask-guided approaches that restrict modifications to a thresholded foreground mask. However, these methods rely on a fixed mask obtained from a single seed, failing to account for the inherent seed variability of diffusion models. Consequently, errors in the mask can lead to significant background inconsistencies.
Candidate Selection for Diffusion Models
Best-of-N is a well-established alignment strategy in inference-time scaling for LLMs [41]. Recent advances have extended inference-time scaling to diffusion models [28, 12], demonstrating the effectiveness of reward models that evaluate the final generated outputs and select the best-of-N result to enhance generation quality. Prior work in T2I generation has explored candidate seed selection [29, 30, 45, 28] and optimization techniques [11, 4], demonstrating that seed choice significantly impacts output quality. However, existing approaches only apply to T2I generation tasks and rely on computationally expensive external models as verifiers [49]. We bridge this gap by introducing a best-of-N selection framework for selecting optimal seeds for I2I editing task.
@@ 3 Preliminaries
Diffusion Models.
Diffusion models [40] involve two processes: a forward process adding noise and a reverse denoising process. Discretized into $T$ timesteps, noise $z_{t}$ is generated with coefficients $\bar{\alpha}_{t}$:
@@
InstructPix2Pix (IP2P).
Recent text-to-image diffusion models [36, 6] are trained on text-conditioned datasets, enabling conditional generation $\epsilon_{\theta}(z_{t},t,C_{T})$ on text embedding $C_{T}$. InstructPix2Pix (IP2P) [1], is a text-conditioned diffusion model fine-tuned on an instruction-based dataset. Built on latent diffusion [36], IP2P learns to modify images by conditioning on both the original image $I$ and an edit instruction $C_{T}$, enabling image-conditional generation $\epsilon_{\theta}(z_{t},t,I,C_{T})$. The strength of the edit can be controlled by the image guidance scale, $s_{I}$ and the text guidance scale $s_{T}$. The final score estimate is then obtained as
@@
Edit Relevance Map.
To evaluate the impact of an edit instruction on an image, we leverage an edit relevance map, first introduced by WYS [31], which estimates the likelihood of each pixel being modified. This map serves as a crucial tool for identifying regions most affected by the edit process.
@@
where $\epsilon\sim\mathcal{N}(0,I)$ is random noise and $\alpha_{t}$ controls the noise level. The IP2P [1] denoising network, $\epsilon_{\theta}$, then predicts noise estimates for both the conditioned and unconditioned cases, $\epsilon_{\theta}(z_{t},t,I,C_{T})$ and $\epsilon_{\theta}(z_{t},t,I,\emptyset)$. The pixel-wise magnitude of their difference provides an estimate of edit relevance $M_{t}=|\epsilon_{\theta}(z_{t},t,I,C_{T})-\epsilon_{\theta}(z_{t},t,I,\emptyset)|$.
Outlier values in $M_{t}$ are clamped using an interquartile range filter and normlized into $[0,1]$.
The same principle applies to rectified flow models [27], which replace diffusion with a velocity field $v_{\theta}$ learned in data space. In this framework, the relevance map is computed as $M_{t}=|v_{\theta}(z_{t},t,I,C_{T})-v_{\theta}(z_{t},t,I,\emptyset)|$, where $v_{\theta}$ guides $z_{t}$ toward the target image. This allows the method to generalize beyond diffusion-based models. For additional details on rectified flow, see Suppl. D.
@@ 4 Method
Figure 3: Overview of the ELECT pipeline (left) and details of Background Inconsistency Score (BIS) computation (right). The left panel illustrates candidate selection via early stopping and BIS evaluation. In the right panel, the top part illustrates the extraction of edit relevance maps for each seed, while the bottom part details the BIS computation process, incorporating crowd-sourced reference masks and background masking to maintain consistency and minimize distortions. The BIS metric compares clean images with the original input to quantify background distortions, ensuring consistent edits with minimal undesired changes.
@@ 4.1 Observations
Image editing models exhibit high output variance across different seeds, with over-editing and distortion levels varying significantly (see Fig. 1). Simply using ground truth (GT) masks, we measured background MSE to identify samples with minimal distortion. We found that selecting the seed with the smallest background MSE effectively reduces unnecessary artifacts and improves instruction following, even without additional training or modulation. These observations are reflected in Fig. 2 where we observed consistent enhancement across multiple metrics.
@@ 4.3 Early-timestep Latent Evaluation for Candidate selecTion (ELECT) Pipeline
13:$i^{*}\leftarrow\arg\min_{i\in\mathbb{S}}S^{\text{BIS}}(i,t_{\text{stop}})$triangleright Select best seed
@@
ELECT is designed to efficiently select the best candidate for image editing while reducing computational costs. As shown in Fig. 3, ELECT evaluates multiple candidates early in the denoising process, eliminating suboptimal ones at an early timestep $t_{\text{stop}}$ before completing inference. ELECT ensures that only the most promising sample is fully denoised, balancing efficiency and accuracy in generative image editing.
Following DDIM [39], $\texttt{Denoise}(z_{t},t,I,C_{T})=\sqrt{\alpha_{t-1}}\hat{z}_{0}+\sqrt{1-\alpha_{t-1}}\cdot\epsilon_{\theta}(z_{t},t,I,C_{T})$ in Algorithm 1. For rectified flow, $\texttt{Denoise}(z_{t},t,I,C_{T})=z_{t}-v_{\theta}(z_{t},t,I,C_{T})\cdot t$ where the time domain is $t\in[0,1]$.
@@ 4.4 ELECT for Instruction Prompt Selection
We found that seed selection improves image editing performance, but its effectiveness plateaus as the number of seeds increases. For out-of-distribution instruction prompts, seed selection alone won’t yield a successful edit. In such cases, modifying the instruction prompt results in improvements (Fig. 4).
To incorporate a Multimodal Large Language Model (MLLM) [34, 49] into the ELECT pipeline, we introduce a new evaluation metric inspired by prior work [21, 22]. This metric classifies edits based on two sub-metrics: Instruction Following and Background Consistency. These sub-metrics independently assess alignment with the instruction and preservation of unedited areas, giving scores of $0,0.5,$ or $1$. If either score is 0, the edit fails, prompting an MLLM to generate alternative instruction prompts while preserving meaning.
@@ 5.1 Setup
Benchmarks and Metrics. Experiments are conducted on PIE-Bench [19], covering 9 editing scenarios with 700 real and synthetic images, and the MagicBrush [50] test set, a manually-annotated real image dataset containing around 560 images. Each dataset provides a source image, edit instruction, and GT foreground mask which is used only for evaluation. Performance is evaluated via CLIPScore [15] for instruction following and PSNR, MSE, SSIM [43], and LPIPS [51] for background consistency. We also report VIEScore [22], a human-aligned metric assessing overall edit quality. For the full implementation detail, refer to Suppl. B.
@@ 5.2 Effect of Seed Selection with ELECT
We evaluate ELECT in terms of both performance and efficiency (denoising steps required) for instruction-guided image editing. As shown in Table 1, multi-seed strategies significantly outperform single-seed evaluation (Vanilla). ELECT achieves the best results, consistently surpassing all baselines across all metrics, particularly when matched to the time complexity of Best of 5 by $S^{\text{BIS}}$. Fig. 7 provides a qualitative comparison across multiple models, comparing Best of 1 (vanilla), Best of 5 with $S^{\text{BIS}}$, and ELECT. Single-seed outputs often exhibit excessive distortion, while selecting a seed with higher background consistency (as in Best of 5) reduces this issue. However, evaluating more seeds within the same computational budget allows ELECT to further minimize unnecessary background modifications. Additional qualitative results are available in Suppl. F.
@@
Figure 6: Quantitative Comparison of MSE between ELECT and Best of N by $S^{\text{BIS}}$. Performance trend (MSE×${}_{\times 10^{4}}$) with respect to the number of function evaluations (NFE), evaluated on two datasets with $t_{\text{stop}}=60$ (Top: PIE-Bench [19], Bottom: MagicBrush [50] test set). Results show consistent performance improvements and enhanced efficiency across various models and datasets. By evaluating more seeds within the same NFE budget, ELECT achieves superior overall performance, surpassing the Pareto front of Best of N by $S^{\text{BIS}}$.
Model Seed Selection Method BC IF VIEScore (Semantic Consistency) ($\uparrow$) Time-complexity (NFE)
@@
5.3 Effect of Instruction Prompt Selection with ELECT
These failure cases are further analyzed in Fig. 8, which presents a comparison of VIEScore metrics before and after prompt selection for samples initially classified as failures following seed selection under ELECT. Simply increasing the number of seed candidates leads to performance saturation or over-optimization, where no further improvements are observed in evaluation metrics. However, across all models, prompt selection effectively overcomes this saturation, resulting in a significant increase in VIEScore. A comprehensive quantitative comparison of all evaluation metrics and more qualitative results are provided in Suppl. F.
Figure 9: ELECT performance variation with respect to stopping timestep ($t_{\text{stop}}$) with fixed number of seeds. This graph shows that performance improves as the stopping denoising step increases, eventually converging around around $t_{\text{stop}}=70$ for most models. In contrast, UltraEdit, as a Rectified Model, exhibits minimal change in noise ratio at very early steps, making it meaningful to select a stopping point after approximately $t_{\text{stop}}=60$.

### v2

@@ Abstract
Despite recent advances in diffusion models, achieving reliable image generation and editing results remains challenging due to the inherent diversity induced by stochastic noise in the sampling process. Particularly, instruction-guided image editing with diffusion models offers user-friendly editing capabilities, yet editing failures, such as background distortion, frequently occur across different attempts. Users often resort to trial and error, adjusting seeds or prompts to achieve satisfactory results, which is inefficient. While seed selection methods exist for Text-to-Image (T2I) generation, they depend on external verifiers, limiting their applicability, and evaluating multiple seeds increases computational complexity, reducing practicality. To address this, we first establish a new multiple-seed-based image editing baseline using background consistency scores, achieving Best-of-N performance without supervision. Building on this, we introduce ELECT (Early-timestep Latent Evaluation for Candidate selecTion), a zero-shot framework that selects reliable seeds by estimating background mismatches at early diffusion timesteps, identfying the seed that retains the background while modifying only the foreground. ELECT ranks seed candidates by a background inconsistency score, filtering unsuitable samples early based on background consistency while fully preserving editability. Beyond standalone seed selection, ELECT integrates into instruction-guided editing pipelines and extends to Multimodal Large-Language Models (MLLMs) for joint seed + prompt selection, further improving results when seed selection alone is insufficient. Experiments show that ELECT reduces computational costs (by 41% on average and up to 61%) while improving background consistency and instruction adherence, achieving around 40% success rates in previously failed cases—without any external supervision or training. Our code is available at https://github.com/Joow0n-Kim/ELECT
@@ 1 Introduction
Auto-regressive models like LLMs face the similar issue, sampling multiple outputs at inference to improve quality [38]. Similarly, in text-to-image generation, search techniques such as Best of N, which select the best result using specific verifiers, have been explored [28]. In instruction-guided image editing, users often generate multiple outputs by varying the random seed and manually selecting the most suitable result. However, this process is computationally expensive, with performance scaling linearly with N, making it impractical for real-time applications. Moreover, there is no established framework for efficiently identifying the optimal seed before full inference, underscoring the need for a more effective selection method.
@@
  • We extend ELECT to joint seed and prompt selection using multimodal large language models (MLLMs) [34] refining out-of-distribution instructions and improving VIEScore by +0.56+0.56+ 0.56 on average.
@@ 2 Related Work
Instruction-Guided Image Editing with Diffusion Models. Unlike caption-based approaches, instruction-guided editing methods [1, 50, 9, 37, 44, 52, 7, 16, 18, 8, 48, 24, 53] take an input image $I$ and a textual command $T$ (e.g., ”add a dog”) to guide modifications. InstructPix2Pix (IP2P) [1] uses GPT-3 [2] and Prompt2Prompt [14] to create a dataset, and trains a denoising network conditioned on edit instructions and the original image. Subsequent studies such as MagicBrush [50], UltraEdit [53], HIVE [52], and HQ-Edit [18] propose fine-tuning techniques for IP2P through improved datasets, verifier models [49], and RLHF [26]. InstructDiffusion (InsDiff) [9] proposes a unified framework across multiple computer vision tasks including segmentation and editing, and MGIE [7] and SmartEdit [16] leverages multi-modal LLMs (MLLMs) to guide and enhance image editing. Although instruction-guided methods outperform previous methods, they tend to overedit images and introduce variability due to sensitivity to initial seeds and instruction phrasing, leading to inconsistent edits across runs. To address this, Watch Your Steps [31], Focus on Your Instruction [10], and ZONE [25] employ mask-guided approaches that restrict modifications to a thresholded foreground mask. However, these methods rely on a fixed mask obtained from a single seed, failing to account for the inherent seed variability of diffusion models. Consequently, errors in the mask can lead to significant background inconsistencies.
Candidate Selection for Diffusion Models. Best-of-N is a well-established alignment strategy in inference-time scaling for LLMs [41]. Recent advances have extended inference-time scaling to diffusion models [28, 12], demonstrating the effectiveness of reward models that evaluate the final generated outputs and select the best-of-N result to enhance generation quality. Prior work in T2I generation has explored candidate seed selection [29, 30, 45, 28] and optimization techniques [11, 4], demonstrating that seed choice significantly impacts output quality. However, existing approaches only apply to T2I generation tasks and rely on computationally expensive external models as verifiers [49]. We bridge this gap by introducing a best-of-N selection framework for selecting optimal seeds for I2I editing task.
@@ 3 Preliminaries
Diffusion Models. Diffusion models [40] involve two processes: a forward process adding noise and a reverse denoising process. Discretized into $T$ timesteps, noise $z_{t}$ is generated with coefficients $\bar{\alpha}_{t}$:
@@
InstructPix2Pix. Recent text-to-image diffusion models [36, 6] are trained on text-conditioned datasets, enabling conditional generation $\epsilon_{\theta}(z_{t},t,C_{T})$ on text embedding $C_{T}$. InstructPix2Pix (IP2P) [1], is a text-conditioned diffusion model fine-tuned on an instruction-based dataset. Built on latent diffusion [36], IP2P learns to modify images by conditioning on both the original image $I$ and an edit instruction $C_{T}$, enabling image-conditional generation $\epsilon_{\theta}(z_{t},t,I,C_{T})$. The strength of the edit can be controlled by the image guidance scale, $s_{I}$ and the text guidance scale $s_{T}$. The final score estimate is then obtained as
@@
Edit Relevance Map. To evaluate the impact of an edit instruction on an image, we leverage an edit relevance map, first introduced by WYS [31], which estimates the likelihood of each pixel being modified. This map serves as a crucial tool for identifying regions most affected by the edit process.
@@
where $\epsilon\sim\mathcal{N}(0,I)$ is random noise and $\alpha_{t}$ controls the noise level. The InstructPix2Pix [1] denoising network, $\epsilon_{\theta}$, then predicts noise estimates for both the conditioned and unconditioned cases, $\epsilon_{\theta}(z_{t},t,I,C_{T})$ and $\epsilon_{\theta}(z_{t},t,I,\emptyset)$. The pixel-wise magnitude of their difference provides an estimate of edit relevance $M_{t}=|\epsilon_{\theta}(z_{t},t,I,C_{T})-\epsilon_{\theta}(z_{t},t,I,\emptyset)|$.
Outlier values in $M_{t}$ are clamped using an interquartile range filter and normlized into [0,1][0,1][ 0 , 1 ].
The same principle applies to rectified flow models [27], which replace diffusion with a velocity field $v_{\theta}$ learned in data space. In this framework, the relevance map is computed as $M_{t}=|v_{\theta}(z_{t},t,I,C_{T})-v_{\theta}(z_{t},t,I,\emptyset)|$, where $v_{\theta}$ guides $z_{t}$ toward the target image. This allows the method to generalize beyond diffusion-based models. For additional details on rectified flow, see App. D.
@@ 4 Method
Figure 2: Overview of the ELECT pipeline (top) and details of Background Inconsistency Score (BIS) computation (bottom). The top panel illustrates candidate selection via early stopping and BIS evaluation. The bottom panel details BIS computation, including crowd-sourced reference masks and background masking, ensuring consistent edits with minimal distortions. The BIS metric compares clean images with the original input to quantify background distortions, ensuring consistent edits while minimal undesired changes.
@@ 4.1 Observations
Image editing models exhibit high output variance across different seeds, with over-editing and distortion levels varying significantly (see LABEL:fig:seed_variance). Simply using ground truth (GT) masks, we measured background MSE to identify samples with minimal distortion. We found that selecting the seed with the smallest background MSE effectively reduces unnecessary artifacts and improves instruction following, even without additional training or modulation. These observations are reflected in Fig. 1 where we observed consistent enhancement across multiple metrics.
@@ 4.3 Early-timestep Latent Evaluation for Candidate selecTion (ELECT) Pipeline
[…]
@@
ELECT is designed to efficiently select the best candidate for image editing while reducing computational costs. As shown in Fig. 2, ELECT evaluates multiple candidates early in the denoising process, eliminating suboptimal ones at an early timestep $t_{\text{stop}}$ before completing inference. ELECT ensures that only the most promising sample is fully denoised, balancing efficiency and accuracy in generative image editing. Following DDIM [39], $\texttt{Denoise}(z_{t},t,I,C_{T})=\sqrt{\alpha_{t-1}}\hat{z}_{0}+\sqrt{1-\alpha_{t-1}}\cdot\epsilon_{\theta}(z_{t},t,I,C_{T})$ in Algorithm 1. For rectified flow, $\texttt{Denoise}(z_{t},t,I,C_{T})=z_{t}-v_{\theta}(z_{t},t,I,C_{T})\cdot\Delta t$ where the time domain is $t\in[0,1]$.
@@ 4.4 ELECT for Instruction Prompt Selection
We found that seed selection improves image editing performance, but its effectiveness plateaus as the number of seeds increases. For out-of-distribution instruction prompts, seed selection alone won’t yield a successful edit. In such cases, modifying the instruction prompt results in improvements (see LABEL:fig:seed_variance).
To incorporate a Multimodal Large Language Model (MLLM) [34, 49] into the ELECT pipeline, we introduce a new evaluation metric inspired by prior work [21, 22]. This metric classifies edits based on two sub-metrics: Instruction Following and Background Consistency. These sub-metrics independently assess alignment with the instruction and preservation of unedited areas, giving scores of 0,0.5,0,0.5,0 , 0.5 , or 111. If either score is 0, the edit fails, prompting an MLLM to generate alternative instruction prompts while preserving meaning. The ELECT pipeline is then applied to select the best prompt early on App. E.
@@ 5.1 Setup
Benchmarks and Metrics. Experiments are conducted on PIE-Bench [19], covering 9 editing scenarios with 700 real and synthetic images, and the MagicBrush [50] test set, a manually-annotated real image dataset containing around 560 images. Each dataset provides a source image, edit instruction, and GT foreground mask which is used only for evaluation. Performance is evaluated via CLIPScore [15] for instruction following and PSNR, MSE, SSIM [43], and LPIPS [51] for background consistency. We also report VIEScore [22], a human-aligned metric assessing overall edit quality. For the full implementation detail, refer to App. B.
@@ 5.2 Effect of Seed Selection with ELECT
We evaluate ELECT in terms of both performance and efficiency (denoising steps required) for instruction-guided image editing. As shown in Tab. 1, multi-seed strategies significantly outperform single-seed evaluation (Vanilla). ELECT achieves the best results, consistently surpassing all baselines across all metrics, particularly when matched to the time complexity of Best of 5 by $S^{\text{BIS}}$. Fig. 5 provides a qualitative comparison across multiple models, comparing Best of 1 (vanilla), Best of 5 with $S^{\text{BIS}}$, and ELECT. Single-seed outputs often exhibit excessive distortion, while selecting a seed with higher background consistency (as in Best of 5) reduces this issue. However, evaluating more seeds within the same computational budget allows ELECT to further minimize unnecessary background modifications. Additional qualitative results are available in App. F.
@@
Model
Seed Selection
Method
BC IF VIEScore (Semantic Consistency) ($\uparrow$)
Time-complexity
@@
5.3 Effect of Instruction Prompt Selection with ELECT
Figure 6: ELECT performance variation with respect to stopping timestep ($t_{\text{stop}}$) with fixed number of seeds. This graph shows that performance improves as the stopping denoising step increases, eventually converging around around $t_{\text{stop}}=70$ for most models. In contrast, UltraEdit, as a Rectified Model, exhibits minimal change in noise ratio at very early steps, making it meaningful to select a stopping point after approximately $t_{\text{stop}}=60$. Full analysis is in the supplementary materials.
These failure cases are further analyzed in Fig. 7, which presents a comparison of VIEScore metrics before and after prompt selection for samples initially classified as failures following seed selection under ELECT. Simply increasing the number of seed candidates leads to performance saturation or over-optimization, where no further improvements are observed in evaluation metrics. However, across all models, prompt selection effectively overcomes this saturation, resulting in a significant increase in VIEScore. A comprehensive quantitative comparison of all evaluation metrics and more qualitative results are provided in App. F.